OpenStudy (anonymous):

What is the sum of the arithmetic sequence 6, 14, 22 …, if there are 26 terms

OpenStudy (vishweshshrimali5):

sum of an AP = n/2 (2a +(n-1)d) Here, n = 26, a = 6, d = 14 - 6 = 8 Now calculate to get the answer

OpenStudy (anonymous):

i got 2754 is that correct?

OpenStudy (vishweshshrimali5):

Let me check

OpenStudy (vishweshshrimali5):

It should be 2756 ?

OpenStudy (anonymous):

ok hold on let me recheck?

OpenStudy (anonymous):

ok i did it wrong i got 2756

OpenStudy (vishweshshrimali5):

Gud work